ANALYSIS OF NEW DATA SETS PERTAINING TO THE WATER MOLECULE

S. YU, J. C. PEARSON, B. J. DROUIN, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109, USA; H. S. P. MÜLLER, S. BRÜNKEN, I. Physikalisches Institut, Universität zu Köln, 50937 Köln, Germany; M. A. MARTIN-DRUMEL, O. PIRALI, D. BALCON, M. VERVLOET, Ligne AILES – Synchrotron SOLEIL, L’Orme des Merisiers, Saint Aubin, 91192 Gif-sur-Yvette, France; AND \underline{L. H. COUDERT}, LISA, CNRS/Universités Paris Est et Paris Diderot, 61 Avenue du Général de Gaulle, 94010 Créteil, France.

Although water is just a simple triatomic molecule, its spectroscopy still remains a challenge due to the ever increasing amount of available data. Two types of data have been recently obtained for water:

• Accurate frequencies have been measured for transitions in the sub millimeter and terahertz domains involving high-lying rovibrational levels up to the first triad. 149 transitions have been measured between 300 GHz and 2 THz and 26 from 2.5 to 2.7 THz. These b-type transitions take place within one of the five first vibrational states.

• Far infrared transitions involving high-lying rovibrational levels have been recorded recently in the 50 to 600 cm^{−1} region using the emission spectrum of a continuous flow of water vapor rovibrationaly excited by an electrodless radio-frequency discharge.^{a} 3793 transitions with ΔK_{a} = 1 within one of the five first vibrational states of the molecule have been assigned so far and involve J-values up to 25 and K_{a}-values up to 15. There remains a large number of unassigned transitions involving either higher lying vibrational states, larger values of ΔK_{a}, or taking place between different vibrational states.

The paper will focus on the results of the analysis of a large data set consisting of already published data^{b} and of the two new data sets. The number of data is equal to 20491 and the bending-rotation theoretical approach^{c} will be used for the energy level calculation.

aPirali and Vervloet, Chem. Phys. Letters 423 (2006) 376.

bCoudert, Wagner, Birk, Baranov, Lafferty, and Flaud, J. Molec. Spec. 251 (2008) 339.

cCoudert, J. Molec. Spec. 181 (1997) 246.